# SEED-Story: Multimodal Long Story Generation with Large Language Model

### v1

@@ Abstract
With the remarkable advancements in image generation and open-form text generation, the creation of interleaved image-text content has become an increasingly intriguing field. Multimodal story generation, characterized by producing narrative texts and vivid images in an interleaved manner, has emerged as a valuable and practical task with broad applications. However, this task poses significant challenges, as it necessitates the comprehension of the complex interplay between texts and images, and the ability to generate long sequences of coherent, contextually relevant texts and visuals. In this work, we propose SEED-Story, a novel method that leverages a Multimodal Large Language Model (MLLM) to generate extended multimodal stories. Our model, built upon the powerful comprehension capability of MLLM, predicts text tokens as well as visual tokens, which are subsequently processed with an adapted visual de-tokenizer to produce images with consistent characters and styles. We further propose multimodal attention sink mechanism to enable the generation of stories with up to 25 sequences (only 10 for training) in a highly efficient autoregressive manner. Additionally, we present a large-scale and high-resolution dataset named StoryStream for training our model and quantitatively evaluating the task of multimodal story generation in various aspects. The models, codes and datasets are released in https://github.com/TencentARC/SEED-Story.
Figure 1: The introduced SEED-Story, powered by MLLM, is capable of generating multimodal long stories from user-provided images and texts as the beginning of the story. The generated story consists of rich and coherent narrative texts, along with images that are consistent in characters and style. The story can span up to 25 multimodal sequences (see Appendix), even though we only use a maximum of 10 sequences during training.
@@ 1 Introduction
Interleaved image-text data is ubiquitous on the internet, characterized by multiple images interspersed with text. In recent years, there has been a surge of interest in generating interleaved image-text content [52, 16, 1, 11, 49], driven by the remarkable advancements in image generation [43, 26, 9, 36, 54] and open-form text generation [53, 48, 59]. This has given rise to Multimodal Story Generation, an intriguing and valuable task that involves the generation of narrative texts and vivid images in an interleaved manner. This task goes beyond traditional text-based storytelling by fusing text and visuals to enrich narratives, creating a more immersive storytelling experience where visuals and text dynamically interact and complement each other.
@@
However, multimodal story generation poses significant challenges due to the complexity of the inputs and the high demands on the quality of the outputs. Firstly, this task necessitates a thorough comprehension of interleaved data, where text is not only abstract and narrative in nature, but also deeply intertwined with complex images. The model must be adept at deciphering the intricate relationships between images and texts to maintain a coherent narrative flow. Secondly, this task requires the generation of not only a plausible text plot, but also visually captivating images that are consistent in characters and styles. The model should be capable of achieving coherence in the generation of both text and visuals, ensuring an engaging storytelling output.
@@
Recently, Multimodal Large Language Models (MLLMs) [23, 60, 39, 5, 28, 58, 26, 22] have showcased powerful comprehension abilities in understanding multimodal data, which makes them ideally suited for interleaved image-text content in multimodal stories. Consequently, we introduce SEED-Story, a novel approach that builds upon the MLLM to harness its comprehension strength, while further equipping it with the capability to generate coherent images align with the narrative texts.
@@
Specifically, following previous work [45, 16], we utilize a pre-trained image tokenizer and de-tokenizer, which can decode realistic images with SD-XL [41] by taking the features of a pre-trained ViT as input. During training, given the interleaved visual and textual data, we adopt the next-word prediction and image feature regression training objectives to regularize multimodal generation. A fixed number of learnable queries are fed into the MLLM, where the output hidden states are trained to reconstruct the ViT features of the target images. To further enhance the consistency of characters and styles in generated images, we propose de-tokenizer adaptation, where the regressed image features from the MLLM are fed into the de-tokenizer for tuning SD-XL. This adaptation allows for better maintenance of coherence in low-level image details from the de-tokenizer, ensuring a more visually consistent storytelling output.
@@
Furthermore, to enable the efficient generation of coherent long stories, we propose a multimodal attention sink mechanism based on window attention [7], which maintains a fixed-size sliding window on the Key-Value (KV) states of the most recent tokens, as well as the beginning of text tokens, images tokens, and the end of image tokens. We empirically find that retaining these tokens will largely mitigate the model’s failure with window attention when the token length surpasses the cache size, allowing our model to generalize to longer sequences than the training sequence length in an efficient manner. Our model with the proposed multimodal attention sink mechanism can generate long stories with up to 30 multimodal sequences, featuring rich text plots and diverse visual scenarios.
@@
Additionally, we introduce a dataset named StoryStream for training and evaluating multimodal story generation. We design an automatic pipeline that leverages MLLMs to obtain a large-scale and high-resolution dataset featuring a sequence of narrative-rich texts and intriguing images, derived from animated videos. StoryStream is four times larger in terms of data volume compared to the existing largest story dataset [27], and it boasts higher image resolution, longer sequence lengths, and more detailed story narratives. We further meticulously design evaluation metrics to assess multimodal story generation, taking into account image style consistency, story engagement, and image-text coherence. The evaluation results demonstrate that our model, SEED-Story, achieves superior performance in these aspects.
@@ 2 Related Work
Story Visualization vs. Multimodal Story Generation
The domain of story visualization, initially pioneered by StoryGAN [25], has witnessed substantial advancements across various models focusing on improving image quality, narrative flow, and character consistency. StoryGAN introduced a framework based on sequential conditional GANs with dual discriminators that enhanced both narrative coherence and visual fidelity. Subsequent developments such as DuCoStoryGAN [32] and VLCStoryGAN [31] have integrated video captioning to better align text descriptions with corresponding visuals. The field has further expanded with the adoption of diffusion models aimed at refining the quality of generated story images. StoryDALL-E [33] utilizes pre-trained models augmented with cross-attention layers to support story progression from an initial image. Innovations like AR-LDM [35] and Story-LDM [42] have introduced auto-regressive diffusion models to create coherent sequences of images, while TaleCrafter [17] has employed LoRA and optimization techniques to ensure consistent characters throughout complex visual narratives. Intelligent Grimm [27] has proposed open-ended story visualization and has developed the auto-regressive model StoryGen, which includes a vision-language context module to further enhance character consistency. In contrast to all above, our work advances beyond story visualization by focusing on multimodal story generation. This approach continues the story with additional text and creates corresponding images, offering a richer, more engaging narrative experience into the story generation process.
@@
In the rapidly evolving domain of large language models (LLMs) [53, 8, 10] and multimodal large language models (MLLMs) [23, 60, 30, 40, 6, 29, 58, 26, 46, 57, 15, 14, 55, 11, 61, 47, 24], recent research, such as StoryGPTV [44], explores using MLLMs for story generation by converting visual features into token embeddings. StoryGPTV requires character and object masks for training its LDM. In contrast, our model, with the visual tokenization and de-tokenization, have a stronger visual comprehension ability, making it trains directly from interleaved text and image data without the need for masks. MM-interleaved [52] design an MMFS module to process interleaved text-image data and achieve multimodal story generation through image features and LLM output. While our approach directly generates images through LLM output tokens. We take MM-interleaved as an important baseline and compare with it.
@@
Datasets Number of Images Resolution Story Length Avg Text Length
Flintstones [18] 122,560 $128\times 128$ 5 86
Pororo [25] 73,665 $128\times 128$ 5 74
StorySalon [27] 159,778 $432\times 803$ 14 106
@@
In the landscape of datasets for visual storytelling, various collections have been developed. The VIST [20] dataset is noteworthy for its use of realistic images, though it struggles with maintaining character consistency across stories. The Pororo [25] and Flintstones [18] datasets, while popular for animation-based story datasets, are hindered by their low resolution and the simplicity of their accompanying texts. Another significant dataset is StorySalon [27], which offers high-resolution images and is large in scale, but it lacks global consistency across a series of images. To address these gaps, we introduce StoryStream, a globally consistent, large-scale, high-resolution animated style dataset with engaging, narrative-rich text for complex storytelling, overcoming the limitations of existing datasets. The analysis is shown in Table 1.
@@ 3.1 Story Generation with Multimodal Large Language Model
The overview of our method is presented in Figure 2. To effectively extend visual stories, our model must comprehend and generate both images and text. Drawing inspiration from recent advancements in generative MLLMs that unify image comprehension and generation [41], we develop a multimodal story generation model. Our model employs a pre-trained Vision Transformer [12] (ViT) as the visual tokenizer and a pre-trained diffusion model as the visual de-tokenizer to decode images by using ViT’s features as inputs. Specifically, visual embeddings from the ViT tokenizer are fed into a learnable module, which then serves as inputs for the U-Net of the pre-trained SD-XL [41]. This process replaces the original text features with visual embeddings. During this stage, the parameters are optimized using open-world text-image pair data as well as story data to enhance the model’s encoding-decoding capability. After this training phase, we expect the visual tokenizer and de-tokenizer modules to preserve as much image information as possible in the feature space.
@@
In our instruction tuning process for story generation, we sample a random-length subset of a story data point for each iteration. The model is tasked with predicting the next image and the next sentence of the story text. Within MLLM, all images are converted into image features using a pre-trained ViT tokenizer. For the target text tokens, we perform next-token prediction and use Cross Entropy loss to train for this discrete target. For the target image features, the model uses a series of learnable queries as inputs and continuously outputs a series of latent embeddings. We then compute the cosine similarity loss between the MLLM’s output and the target image features. During this stage, we fine-tune the SEED-Story model using a LoRA [19] module.
@@
After instruction tuning, the SEED-Story MLLM effectively produces story images with correct semantics but lacks style consistency and details. We attribute this issue to the misalignment between the latent space of the MLLM output and the image features. To address this, we perform de-tokenizer adaptation for style and texture alignment. In this stage, only the SD-XL image de-tokenizer is trained. Conditioned on the MLLM output embeddings, SD-XL is expected to generate images that are pixel-level aligned with the ground truth. The separate training of the de-tokenizer offers two key advantages. First, it avoids optimization conflicts between the LLM and the de-tokenizer. Second, it conserves memory, making the process executable on GPUs with limited memory.
@@ 3.2 Long Story Generation with Multimodal Attention Sink
A simplistic solution for this is to use a sliding window technique, depicted in Figure 3 right (b). However, this method disrupts the token relationships in the Key-Value (KV) cache, resulting in subpar generative outcomes, as demonstrated by StreamingLLM [56]. To overcome this, StreamingLLM introduces an attention sink mechanism that preserves the initial tokens, thus allowing for efficient processing of lengthy generations without quality compromise. While effective in language models, its efficacy diminishes in multimodal contexts, as shown in Figure 3 right (c).
To enhance long multimodal generation, we revisit the attention maps of MLLMs. After conducting numerous experiments across various models and cases, we analyze the attention maps across different layers and heads. Our analysis reveals that most queries predominantly focus on four types of tokens: (1) starting tokens, (2) punctuation tokens, (3) beginning-of-image (BoI) tokens, and (4) end-of-image (EoI) tokens. Unlike language-only models, MLLMs place considerable attention on specific image tokens, particularly those near the BoI and EoI, as illustrated in Figure 3 left.
Building on these insights, we propose a new mechanism for extended generation in MLLMs, termed the multimodal attention sink. During generation, we consistently retain the starting tokens and the image tokens adjacent to the BoI and EoI. Although punctuation tokens receive high attention values, their latent value norms are minimal, contributing insignificantly to the final output, so we do not keep them, as noted by [13]. Our proposed mechanism enables our model to generate high-quality images while maintaining a low computational footprint.
@@ 4.1 Dataset Construction
An ideal source for creating a multimodal story generation dataset is cartoon series, which inherently contain rich plots and consistent character portrayals. We selected three cartoon series to construct our dataset and we present the Curious George in the main body of our paper. The process begins with collecting various series, from which we extract keyframes and their associated subtitles [21]. Each keyframe is then processed by GPT-4V [34] or Qwen-VL [6] to generate a detailed image description. These elements—keyframe, subtitle, and description—are compiled into a single group. We aggregate 30 such groups and input them into GPT-4, supplemented with background information about the cartoon series. Following our instructions, GPT-4 generates high-quality narrative texts suitable for training story generation models.

### v2

@@ Abstract
With the remarkable advancements in image generation and open-form text generation, the creation of interleaved image-text content has become an increasingly intriguing field. Multimodal story generation, characterized by producing narrative texts and vivid images in an interleaved manner, has emerged as a valuable and practical task with broad applications. However, this task poses significant challenges, as it necessitates the comprehension of the complex interplay between texts and images, and the ability to generate long sequences of coherent, contextually relevant texts and visuals. In this work, we propose SEED-Story, a novel method that leverages a Multimodal Large Language Model (MLLM) to generate extended multimodal stories. Our model, built upon the powerful comprehension capability of MLLM, predicts text tokens as well as visual tokens, which are subsequently processed with an adapted visual de-tokenizer to produce images with consistent characters and styles. We further propose multimodal attention sink mechanism to enable the generation of stories with up to 25 sequences (only 10 for training) in a highly efficient autoregressive manner. Additionally, we present a large-scale and high-resolution dataset named StoryStream for training our model and quantitatively evaluating the task of multimodal story generation in various aspects. All models, training and inference codes are released in https://github.com/TencentARC/SEED-Story.
Figure 1: The introduced SEED-Story, powered by MLLM, is capable of generating multimodal long stories based on user-provided images and texts as the story’s beginning. The generated story features rich and coherent narrative texts, accompanied by images that maintain consistency in characters and style. The story can span up to 25 multimodal sequences (see Appendix), even though we only use a maximum of 10 sequences for training.
@@ 1 Introduction
Interleaved image-text data is ubiquitous on the internet, characterized by multiple images interspersed with text. In recent years, there has been a surge of interest in generating interleaved image-text content Tian et al. (2024); Ge et al. (2024); Aiello et al. (2023); Dong et al. (2023); Team (2024), driven by the remarkable advancements in image generation Rombach et al. (2022); Lin et al. (2023); Chen et al. (2023); Patashnik et al. (2021); Wang et al. (2023) and open-form text generation Touvron et al. (2023); Taori et al. (2023); Zheng et al. (2023). This has given rise to Multimodal Story Generation, an intriguing and valuable task that involves the generation of a sequence of consecutive storylines along with their corresponding vivid images in an interleaved manner, similar to that of a serialized comic.
@@
Different from Personalized Story Visualization Ruiz et al. (2023); Avrahami et al. (2024); Tewel et al. (2024), which aims to generate consistent images based on the provided captions following the pattern of text-to-image generation, multimodal story generation poses a more significant challenge due to the complexity of the inputs and the high demands of the outputs. Firstly, this task necessitates a thorough comprehension of interleaved data, where text is not only abstract and narrative in nature, but also deeply intertwined with complex images. The model must be adept at deciphering the intricate relationships between images and texts to maintain a coherent narrative flow. Secondly, this task requires the generation of not only a plausible text plot, but also visually captivating images that are consistent in characters and styles. The model should be capable of achieving coherence in the generation of both text and visuals, ensuring an engaging storytelling output.
@@
Recently, Multimodal Large Language Models (MLLMs) Li et al. (2023); Zhu et al. (2023a); Peng et al. (2023a); Bai et al. (2023a); Liu et al. (2023c); Zhang et al. (2023); Lin et al. (2023); Laurençon et al. (2024) have showcased powerful comprehension abilities in understanding multimodal data, which makes them ideally suited for interleaved image-text content in multimodal stories. Consequently, we introduce SEED-Story, as shown in Figure 1, a novel approach that builds upon the MLLM to harness its comprehension strength, while further equipping it with the capability to generate coherent images align with the narrative texts.
@@
Specifically, following previous work Sun et al. (2023a); Ge et al. (2024), we utilize a pre-trained image tokenizer and de-tokenizer, which can decode realistic images with SD-XL Podell et al. (2023) by taking the features of a pre-trained ViT as input. During training, given the interleaved visual and textual data, we adopt the next-word prediction and image feature regression training objectives to regularize multimodal generation. A fixed number of learnable queries are fed into the MLLM, where the output hidden states are trained to reconstruct the ViT features of the target images. To further enhance the consistency of characters and styles in generated images, we propose de-tokenizer adaptation, where the regressed image features from the MLLM are fed into the de-tokenizer for tuning SD-XL. This adaptation allows for better maintenance of coherence in low-level image details from the de-tokenizer, ensuring a more visually consistent storytelling output.
@@
Furthermore, to enable the efficient generation of coherent long stories, we propose a multimodal attention sink mechanism based on window attention Beltagy et al. (2020), which maintains a fixed-size sliding window on the Key-Value (KV) states of the most recent tokens, as well as the beginning of text tokens, images tokens, and the end of image tokens. We empirically find that retaining these tokens will largely mitigate the model’s failure with window attention when the token length surpasses the cache size, allowing our model to generalize to longer sequences than the training sequence length in an efficient manner. Our model with the proposed multimodal attention sink mechanism can generate long stories with up to 30 multimodal sequences, featuring rich text plots and diverse visual scenarios.
@@
Additionally, we introduce a dataset named StoryStream for training and evaluating multimodal story generation. We design an automatic pipeline that leverages MLLMs to obtain a large-scale and high-resolution dataset featuring a sequence of narrative-rich texts and intriguing images, derived from animated videos. StoryStream is four times larger in terms of data volume compared to the existing largest story dataset Liu et al. (2023a), and it boasts higher image resolution, longer sequence lengths, and more detailed story narratives. We further meticulously design evaluation metrics to assess multimodal story generation, taking into account image style consistency, story engagement, and image-text coherence. The evaluation results demonstrate that our model, SEED-Story, achieves superior performance in these aspects.
@@ 2 Related Work
Figure 2: Comparison between personalized story visualization and multimodal story generation. Contents in grey boxes refer to user’s input. In the former, a sequence of captions are given (referred to as a “story”) for consistent text-to-image generation. By contrast, multimodal story generation involves creating a sequence of consecutive storylines along with their corresponding images.
Personalized Story Visualization v.s. Multimodal Story Generation
Personalized Story Visualization, such as Li et al. (2019); Maharana et al. (2021); Maharana & Bansal (2021); Maharana et al. (2022); Pan et al. (2024); Rahman et al. (2023); Gong et al. (2023); Liu et al. (2023a); Ruiz et al. (2023); Avrahami et al. (2024); Tewel et al. (2024), aims to generate images depicting specified characters engaged in various actions or within different scenes, based on the provided captions (so-called "story") as shown in Fig. 2, which follows the pattern of text-to-image generation. For example, StoryDALL-E Maharana et al. (2022) utilizes pre-trained models augmented with cross-attention layers to support story progression from an initial image. Innovations like AR-LDM Pan et al. (2024) and Story-LDM Rahman et al. (2023) have introduced auto-regressive diffusion models to create coherent sequences of images, while TaleCrafter Gong et al. (2023) has employed LoRA and optimization techniques to ensure consistent characters throughout complex visual narratives.
Multimodal Story Generation (Shen & Elhoseiny (2023)) aims to generate a sequence of consecutive storylines along with their corresponding images, similar to that of a serialized comic. To achieve this, a model must be capable of predicting reasonable story developments and generating corresponding illustrations, by incorporating the previous results as context in an auto-regressive manner. As shown in Figure 2, the generated story images exhibit rapidly changing backgrounds and characters, different from personalized story visualization where the main character consistently appears in the images.
The task of multimodal story generation presents a more substantial challenge, and we follow previous research Shen & Elhoseiny (2023) to adopt a closed-set setting. We believe the ultimate goal of multimodal story generation should be to generate highly diverse scenarios while also generalizing to unseen characters, which will be explored in our future work.
@@
In the rapidly evolving domain of large language models (LLMs) Touvron et al. (2023); Brown et al. (2020); Chowdhery et al. (2022) and multimodal large language models (MLLMs) Li et al. (2023); Zhu et al. (2023a); Liu et al. (2023d); Peng et al. (2023b); Bai et al. (2023b); Liu et al. (2023b); Zhang et al. (2023); Lin et al. (2023); Sun et al. (2023c); Yu et al. (2023); Ge et al. (2023c; b); Wu et al. (2023); Dong et al. (2023); Zhu et al. (2023b); Sun et al. (2023b); Li et al. (2024), recent work StoryGPTV Shen & Elhoseiny (2023) explores using MLLMs for story generation by converting visual features into token embeddings for image generation, but requires aditional character and object masks for training. MM-interleaved Tian et al. (2024) designs a multi-scale and multi-image feature synchronizer module (MMFS) to process interleaved text-image data and generates next image conditioned on the previous context features from LLM, which makes it difficult to generate long multimodal stories due to the complex multi-scale attention mechanism.
@@
Datasets Number of Resolution Story Avg Text
Images Length Length
Flintstones Gupta et al. (2018) 122,560 $128\times 128$ 5 86
Pororo Li et al. (2019) 73,665 $128\times 128$ 5 74
StorySalon Liu et al. (2023a) 159,778 $432\times 803$ 14 106
@@
In the landscape of datasets for visual storytelling, various collections have been developed as shown in Table 1. The VIST Huang et al. (2016) dataset is noteworthy for its use of realistic images, though it struggles with maintaining character consistency across stories. Pororo Li et al. (2019) and Flintstones Gupta et al. (2018) datasets, while popular for animation-based story datasets, are hindered by their low resolution and the simplicity of their accompanying texts. Another significant dataset is StorySalon Liu et al. (2023a), which offers high-resolution images and is large in scale, but it lacks global consistency across images. To address these gaps, we introduce StoryStream, a globally consistent, large-scale, high-resolution animated style dataset with engaging, narrative-rich text for complex storytelling, overcoming the limitations of existing datasets.
@@ 3.1 Story Generation with Multimodal Large Language Model
The overview of our method is presented in Figure 3. To effectively extend visual stories, our model must comprehend and generate both images and text. Drawing inspiration from recent advancements in generative MLLMs that unify image comprehension and generation Podell et al. (2023), we develop a multimodal story generation model. Our model employs a pre-trained Vision Transformer Dosovitskiy et al. (2020) (ViT) as the visual tokenizer and a pre-trained diffusion model as the visual de-tokenizer to decode images by using ViT’s features as inputs. Specifically, visual embeddings from the ViT tokenizer are fed into a learnable module, which then serves as inputs for the U-Net of the pre-trained SD-XL Podell et al. (2023). This process replaces the original text features with visual embeddings. During this stage, the parameters are optimized using open-world text-image pair data as well as story data to enhance the model’s encoding-decoding capability. After this training phase, we expect the visual tokenizer and de-tokenizer modules to preserve as much image information as possible in the feature space.
@@
In our instruction tuning process for story generation, we sample a random-length subset of a story data point for each iteration. The model is tasked with predicting the next image and the next sentence of the story text. Within MLLM, all images are converted into image features using a pre-trained ViT tokenizer. For the target text tokens, we perform next-token prediction and use Cross Entropy loss to train for this discrete target. For target image features, the model uses a series of learnable queries as inputs and continuously outputs a series of latents. We then compute the cosine similarity loss between the MLLM’s output and the target image features. During this stage, we fine-tune the SEED-Story model using a LoRA Hu et al. (2021) module.
@@
After instruction tuning, the SEED-Story MLLM effectively produces story images with correct semantics but lacks style consistency and details. We attribute this issue to the misalignment between the latent space of the MLLM output and the image features. To address this, we perform de-tokenizer adaptation for style and texture alignment. In this stage, only the SD-XL image de-tokenizer is trained. Conditioned on the MLLM output embeddings, SD-XL is expected to generate images that are pixel-level aligned with the ground truth. The separate training of the de-tokenizer offers two key advantages. First, it avoids optimization conflicts between the LLM and the de-tokenizer. Second, it conserves memory, making the process executable on GPUs with limited memory. Please see more analysis in Section B of our appendix.
@@ 3.2 Long Story Generation with Multimodal Attention Sink
A simplistic solution for this is to use a sliding window technique, depicted in Figure 4 right (b). However, this method disrupts the token relationships in the Key-Value (KV) cache, resulting in subpar generative outcomes, as demonstrated by StreamingLLM Xiao et al. (2023). To overcome this, StreamingLLM introduces an attention sink mechanism that preserves the initial tokens, thus allowing for efficient processing of lengthy generations without quality compromise. While effective in language models, its efficacy diminishes in multimodal contexts, as shown in Figure 4 right (c).
To enhance long multimodal generation, we revisit the attention maps of MLLMs. After conducting numerous experiments (see Section D of the appendix for more details) across various models and cases, we analyze the attention maps across different layers and heads. Our analysis reveals that most queries predominantly focus on four types of tokens: (1) starting tokens, (2) punctuation tokens, (3) beginning-of-image (BoI) tokens, and (4) end-of-image (EoI) tokens. Unlike language-only models, MLLMs place considerable attention on specific image tokens, particularly those near the BoI and EoI, as illustrated in Figure 4 left.
Building on these insights, we propose a new mechanism for extended generation in MLLMs, termed the multimodal attention sink. During generation, we consistently retain the starting tokens and the image tokens adjacent to the BoI and EoI. Although punctuation tokens receive high attention values, their latent value norms are minimal, contributing insignificantly to the final output, so we do not keep them, as noted by Ge et al. (2023a). Our proposed mechanism enables our model to generate high-quality images while maintaining a low computational footprint.
@@ 4.1 Dataset Construction
An ideal source for creating a multimodal story generation dataset is cartoon series, which inherently contain rich plots and consistent character portrayals. We selected three cartoon series to construct our dataset and we present the Curious George in the main body of our paper. The process begins with collecting various series, from which we extract keyframes and their associated subtitles Kilian et al. (2023). Each keyframe is then processed by GPT-4V OpenAI (2023) or Qwen-VL Bai et al. (2023b) to generate a detailed image description. These elements—keyframe, subtitle, and description—are compiled into a single group. We aggregate 30 such groups and input them into GPT-4, supplemented with background information about the cartoon series. Following our instructions, GPT-4 generates high-quality narrative texts suitable for training story generation models.